# The numeraire e-variable

### 29 Feb 2024 v1

@@ 5 Examples
We now turn to examples. In these examples we consider a single random observation $Z$ in a measurable space ${\mathcal{Z}}$. For concreteness we take $(\Omega,{\mathcal{F}})$ to be ${\mathcal{Z}}$ with its $\sigma$-algebra, and $Z$ the canonical random variable. We will always have ${\mathbb{Q}}\ll{\mathcal{P}}$ in these examples. We start with a simple parametric example, before moving on to more sophisticated nonparametric examples where there exists no reference measure, in order to display the power and generality of the above results. Except for the first example, the others cannot be handled by past work like Wasserman et al. (2020); Grünwald et al. (2024); Lardy et al. (2023).
@@ 5.1 A parametric null
We start with a simple parametric example from Lardy et al. (2023) involving a real-valued observation $Z$ in ${\mathcal{Z}}={\mathbb{R}}$, and then generalize our observations at the end of this subsection. The null hypothesis ${\mathcal{P}}$ consists of two unit-variance Gaussians ${\mathbb{P}}_{1}$ and ${\mathbb{P}}_{2}$ with mean $+1$ and $-1$, and the alternative hypothesis ${\mathbb{Q}}$ is the standard Cauchy distribution. We claim that the RIPr is *${\mathbb{P}}^{*}=({\mathbb{P}}_{1}+{\mathbb{P}}_{2})/2$. This probability measure is equivalent to ${\mathbb{Q}}$ and belongs to ${\mathcal{P}}^{\circ\circ}$, which always contains the convex hull of ${\mathcal{P}}$. Furthermore, writing $p_{1},p_{2},p^{*},q$ for the Lebesgue densities, we have $p_{2}(z)=p_{1}(-z)$ and $q(z)=q(-z)$ and hence
@@
Since $p_{1}+p_{2}=2p^{*}$ we also have ${\mathbb{E}}_{\mathbb{Q}}[p_{1}(Z)/p^{*}(Z)]+{\mathbb{E}}_{\mathbb{Q}}[p_{2}(Z)/p^{*}(Z)]=2$, and conclude that
${\mathbb{E}}_{\mathbb{Q}}\left[\frac{p_{i}(Z)}{p^{*}(Z)}\right]=1,\quad i=1,2.$
Thus (4.3) of Theorem 4.6 is satisfied, and ${\mathbb{P}}^{*}$ is the RIPr. The numeraire is
@@
It is worth remarking that the universal inference e-variable (Wasserman et al., 2020) would be given by $q(Z)/\max(p_{1}(Z),p_{2}(Z))$, which is clearly smaller than $X^{*}$. Note also that we could have taken ${\mathbb{Q}}$ to be any symmetric distribution with a strictly positive density. More generally, we could let ${\mathbb{Q}}$ be symmetric with a density $q$ that is not necessarily strictly positive. In this case the RIPr would be a sub-probability measure with density $p^{*}=\frac{1}{2}(p_{1}+p_{2})\bm{1}_{\{q>0\}}$, and the numeraire would still be given by (5.1).
@@ 5.2 Bounded mean
This example is a variant of one studied in Waudby-Smith and Ramdas (2023). Let ${\mathcal{Z}}=[0,1]$ so that the random variable $Z$ takes values in the unit interval. Fix $\mu\in(0,\frac{1}{2})$ and consider the null hypothesis that the mean of $Z$ is at most $\mu$,
${\mathcal{P}}=\{{\mathbb{P}}\in M_{1}\colon{\mathbb{E}}_{\mathbb{P}}[Z]\leq\mu\}.$
There is no single dominating measure for ${\mathcal{P}}$ since it contains the uncountable non-dominated family $\{\delta_{z}\colon z\in[0,\mu]\}$. However, ${\mathcal{P}}$ is generated by ${\mathcal{E}}_{0}=\{Z/\mu\}$ in the sense of Corollary 4.2, so our strategy will be to locate a candidate numeraire and then apply the corollary to verify that the candidate is, in fact, the numeraire. The alternative hypothesis ${\mathbb{Q}}$ is the uniform distribution on $[0,1]$, and we have ${\mathbb{Q}}\ll{\mathcal{P}}$ for the simple reason that ${\mathbb{P}}(A)=0$ for all ${\mathbb{P}}\in{\mathcal{P}}$ implies that $A$ must actually be empty.
To find a candidate numeraire, we observe that two natural e-variables are $Z/\mu$ and the constant one. All convex combinations of these are also e-variables; equivalently, $1+\lambda(Z-\mu)$ is an e-variable for each $\lambda\in[0,\mu^{-1}]$. We now look for a log-optimal e-variable in this class by directly maximizing $f(\lambda)={\mathbb{E}}_{\mathbb{Q}}[\log(1+\lambda(Z-\mu))]$ over $\lambda\in[0,\mu^{-1}]$. This is a strictly concave function whose derivative $f^{\prime}(\lambda)={\mathbb{E}}_{\mathbb{Q}}[(Z-\mu)/(1+\lambda(Z-\mu))]$ satisfies $f^{\prime}(0)=\frac{1}{2}-\mu>0$ and $f^{\prime}(\mu^{-1})={\mathbb{E}}_{\mathbb{Q}}[\mu-\mu^{2}/Z]=-\infty$. Thus there is a unique interior maximizer $\lambda^{*}\in(0,\mu^{-1})$ which is characterized by the first-order condition
${\mathbb{E}}_{\mathbb{Q}}\left[\frac{Z-\mu}{1+\lambda^{*}(Z-\mu)}\right]=0.$ (5.2)
Since ${\mathbb{Q}}$ is the standard uniform distribution we can be more explicit. Nothing changes if we first multiply both sides by $\lambda^{*}$, and then the left-hand side becomes
@@
${\mathbb{E}}_{\mathbb{Q}}\left[\frac{X}{X^{*}}\right]=1+(\lambda-\lambda^{*}){\mathbb{E}}_{\mathbb{Q}}\left[\frac{Z-\mu}{1+\lambda^{*}(Z-\mu)}\right]=1.$ (5.3)
@@
(Recall that $F_{+}$ denotes the set of all $[0,\infty]$-valued measurable functions.) It is clear that every such $X$ is an e-variable, but it is less obvious that every e-variable is of this form. Nonetheless, equipped with this knowledge the calculation in (5.3) would directly yield the numeraire property of $X^{*}$. The usefulness of Corollary 4.2 is that a complete description of ${\mathcal{E}}$ is not required; it is enough to check (5.3) for $X=1$ and for $X$ in the generating set ${\mathcal{E}}_{0}$.
@@ 5.3 Sub-Gaussian with nonpositive mean
${\mathcal{P}}=\left\{{\mathbb{P}}\in M_{1}\colon{\mathbb{E}}_{\mathbb{P}}[e^{\lambda Z-\lambda^{2}/2}]\leq 1\text{ for all }\lambda\in[0,\infty)\right\}.$
The ‘one-sided’ restriction $\lambda\in[0,\infty)$ implies that $Z$ has nonpositive (potentially nonzero and even infinite) mean under any ${\mathbb{P}}\in{\mathcal{P}}$. Indeed, monotone convergence and the definition of ${\mathcal{P}}$ yield ${\mathbb{E}}_{\mathbb{P}}[Z]=\lim_{\lambda\downarrow 0}{\mathbb{E}}_{\mathbb{P}}[(e^{\lambda Z}-1)/\lambda]\leq\lim_{\lambda\downarrow 0}(e^{\lambda^{2}/2}-1)/\lambda=0$. As in the previous example, ${\mathcal{P}}$ does not admit any dominating measure. It is generated by the family ${\mathcal{E}}_{0}=\{e^{\lambda Z-\lambda^{2}/2}\colon\lambda\in[0,\infty)\}$, so we will again look for a candidate numeraire and verify it using Corollary 4.2. We let the alternative hypothesis ${\mathbb{Q}}$ be normal with mean $\mu>0$ and unit variance. As in the previous example, and for the same reason, we have ${\mathbb{Q}}\ll{\mathcal{P}}$.
To find a candidate numeraire we maximize ${\mathbb{E}}_{\mathbb{Q}}[\log X]$ over $X\in{\mathcal{E}}_{0}$. That is, we maximize ${\mathbb{E}}_{\mathbb{Q}}[\lambda Z-\lambda^{2}/2]=\lambda\mu-\lambda^{2}/2$ over $\lambda\in[0,\infty)$. The maximizer is $\lambda^{*}=\mu$, which yields the candidate
@@
${\mathbb{E}}_{\mathbb{Q}}\left[\frac{e^{\lambda Z-\lambda^{2}/2}}{e^{\mu Z-\mu^{2}/2}}\right]={\mathbb{E}}_{\mathbb{Q}}\left[e^{(\lambda-\mu)(Z-\mu)-(\lambda-\mu)^{2}/2}\right]=1.$ (5.4)
@@
${\mathbb{E}}_{\mathbb{Q}}\left[\frac{X}{X^{*}}\right]\leq\int_{[0,\infty)}{\mathbb{E}}_{\mathbb{Q}}\left[\frac{e^{\lambda Z-\lambda^{2}/2}}{e^{\mu Z-\mu^{2}/2}}\right]\pi(d\lambda)=1.$
@@ 5.4 Symmetric distributions
${\mathcal{P}}=\left\{{\mathbb{P}}\in M_{1}\colon Z\text{ and }-Z\text{ have the same distribution under }{\mathbb{P}}\right\},$
which is again a non-dominated family. We also fix an alternative hypothesis ${\mathbb{Q}}$ that admits a Lebesgue density $q$. It is natural to conjecture that the RIPr is given by the symmetrization $\widetilde{\mathbb{P}}$ of ${\mathbb{Q}}$, whose density is $\widetilde{p}(z)=(q(z)+q(-z))/2$. However, this cannot quite be true in general because $\widetilde{\mathbb{P}}$ need not be equivalent to ${\mathbb{Q}}$. Instead, we claim that the RIPr is the measure ${\mathbb{P}}^{*}$ with density
@@
This is the absolutely continuous part of $\widetilde{\mathbb{P}}$ with respect to ${\mathbb{Q}}$. It is a probability measure if ${\mathbb{Q}}$ has symmetric support, and otherwise a proper sub-probability measure. Note that $\widetilde{\mathbb{P}}$ belongs to ${\mathcal{P}}$, which in particular shows that ${\mathbb{Q}}\ll{\mathcal{P}}$ since ${\mathbb{Q}}\ll\widetilde{\mathbb{P}}$. (Alternatively, we again have that ${\mathbb{P}}(A)=0$ for all ${\mathbb{P}}\in{\mathcal{P}}$ implies that $A$ is empty, which also yields ${\mathbb{Q}}\ll{\mathcal{P}}$.)
To check that ${\mathbb{P}}^{*}$ is the RIPr we will show that the implied candidate numeraire is, in fact, the numeraire. It is given by
$X^{*}=\frac{d{\mathbb{Q}}}{d{\mathbb{P}}^{*}}=\frac{2q(Z)}{q(Z)+q(-Z)}.$
@@
where we recall that a function $\phi$ is odd if $\phi(z)+\phi(-z)=0$ for all $z\in{\mathbb{R}}$. Indeed, any $X$ of this form satisfies ${\mathbb{E}}_{\mathbb{P}}[X]\leq 1+{\mathbb{E}}_{\mathbb{P}}[\phi(Z)]=1$, where the symmetry of $Z$ and the oddness of $\phi$ were used to get ${\mathbb{E}}_{\mathbb{P}}[\phi(Z)]={\mathbb{E}}_{\mathbb{P}}[\phi(-Z)]=-{\mathbb{E}}_{\mathbb{P}}[\phi(Z)]$ and hence ${\mathbb{E}}_{\mathbb{P}}[\phi(Z)]=0$. Conversely, for any e-variable $X=f(Z)$ we may write $X=\frac{1}{2}(f(Z)+f(-Z))+\phi(Z)$ where $\phi(z)=\frac{1}{2}(f(z)-f(-z))$ is the odd part of $f(z)$. For any $z\in{\mathbb{R}}$ we use the symmetric distribution ${\mathbb{P}}=\frac{1}{2}(\delta_{z}+\delta_{-z})$ and the fact that $X$ is an e-variable to get $\frac{1}{2}(f(z)+f(-z))={\mathbb{E}}_{\mathbb{P}}[f(Z)]\leq 1$. Hence $X\leq 1+\phi(Z)$ as required. We could also have reached this conclusion by using that every admissible e-variable is of the form $1+\phi(Z)$ for some odd function $\phi$ as shown by Ramdas et al. (2020).
We can now verify the numeraire property. First, $X^{*}$ is ${\mathbb{Q}}$-almost surely strictly positive and finite, and it is an e-variable because it can be written as $X^{*}=1+\phi^{*}(Z)$ where
@@
${\mathbb{E}}_{\mathbb{Q}}\left[\frac{X}{X^{*}}\right]\leq{\mathbb{E}}_{\mathbb{Q}}\left[(1+\phi(Z))\frac{q(Z)+q(-Z)}{2q(Z)}\right]={\mathbb{E}}_{\widetilde{\mathbb{P}}}\left[(1+\phi(Z))\mathbf{1}_{\{q(Z)>0\}}\right]\leq{\mathbb{E}}_{\widetilde{\mathbb{P}}}[1+\phi(Z)]=1.$
This confirms the numeraire property. Finally, let us remark that it is not really necessary that ${\mathbb{Q}}$ admit a density. The symmetrization of ${\mathbb{Q}}$ is still well-defined as ~$\widetilde{\mathbb{P}}=\frac{1}{2}({\mathbb{Q}}+\widetilde{\mathbb{Q}})$ where $\widetilde{\mathbb{Q}}$ is the distribution of $-Z$ under ${\mathbb{Q}}$, or equivalently, the pushforward of ${\mathbb{Q}}$ under the reflection map $z\mapsto-z$. The RIPr is then the absolutely continuous part *=~${\mathbb{P}}^{*}=\widetilde{\mathbb{P}}^{a}$, and the numeraire is any nonnegative version of the Radon–Nikodym derivative $d{\mathbb{Q}}/d{\mathbb{P}}^{*}$ such that $d{\mathbb{Q}}/d{\mathbb{P}}^{*}-1$ is odd.

### 22 Mar 2024 v2

@@ 5 Examples
We now turn to examples. In these examples we consider a single random observation $Z$ in a measurable space ${\mathcal{Z}}$. For concreteness we take $(\Omega,{\mathcal{F}})$ to be ${\mathcal{Z}}$ with its $\sigma$-algebra, and $Z$ the canonical random variable. We will always have ${\mathsf{Q}}\ll{\mathcal{P}}$. We start with a simple parametric example, before moving on to more sophisticated nonparametric examples where there exists no reference measure, in order to display the power and generality of our theory. We end with another parametric example. Except for the first and the last example, past work like Wasserman et al. (2020); Grünwald et al. (2024); Lardy et al. (2023) cannot handle the situations discussed here.
@@ 5.1 A parametric null
We start with a simple parametric example from Lardy et al. (2023) involving a real-valued observation $Z$ in ${\mathcal{Z}}={\mathbb{R}}$, and then generalize our observations at the end of this subsection. The null hypothesis ${\mathcal{P}}$ consists of two unit-variance Gaussians ${\mathsf{P}}_{1}$ and ${\mathsf{P}}_{2}$ with mean $+1$ and $-1$, and the alternative hypothesis ${\mathsf{Q}}$ is the standard Cauchy distribution. We claim that the RIPr is ${\mathsf{P}}^{*}=({\mathsf{P}}_{1}+{\mathsf{P}}_{2})/2$. This probability measure is equivalent to ${\mathsf{Q}}$ and belongs to ${\mathcal{P}}^{\circ\circ}$, which always contains the convex hull of ${\mathcal{P}}$. Furthermore, writing $p_{1},p_{2},p^{*},q$ for the Lebesgue densities, we have $p_{2}(z)=p_{1}(-z)$ and $q(z)=q(-z)$ and hence
@@
Since $p_{1}+p_{2}=2p^{*}$ we also have ${\mathbb{E}}_{\mathsf{Q}}[p_{1}(Z)/p^{*}(Z)]+{\mathbb{E}}_{\mathsf{Q}}[p_{2}(Z)/p^{*}(Z)]=2$, and conclude that
${\mathbb{E}}_{\mathsf{Q}}\left[\frac{p_{i}(Z)}{p^{*}(Z)}\right]=1,\quad i=1,2.$
Thus (4.3) of Theorem 4.7 is satisfied, and ${\mathsf{P}}^{*}$ is the RIPr. The numeraire is
@@
It is worth remarking that the universal inference e-variable (Wasserman et al., 2020) would be given by $q(Z)/\max\{p_{1}(Z),p_{2}(Z)\}$, which is clearly smaller than $X^{*}$. Note also that we could have taken ${\mathsf{Q}}$ to be any symmetric distribution with a strictly positive density. More generally, we could let ${\mathsf{Q}}$ be symmetric with a density $q$ that is not necessarily strictly positive. In this case the RIPr would be a sub-probability measure with density $p^{*}=\frac{1}{2}(p_{1}+p_{2})\bm{1}_{\{q>0\}}$, and the numeraire would still be given by (5.1).
@@ 5.2 Bounded mean
This example is a variant of one studied in Waudby-Smith and Ramdas (2023), with the corresponding duality theory derived explicitly in Honda and Takemura (2010), and generalized to other classes of distributions in Agrawal et al. (2020). Let ${\mathcal{Z}}=[0,1]$ so that the random variable $Z$ takes values in the unit interval. Fix $\mu\in(0,\frac{1}{2})$ and consider the null hypothesis that the mean of $Z$ is at most $\mu$,
${\mathcal{P}}=\{{\mathsf{P}}\in M_{1}\colon{\mathbb{E}}_{\mathsf{P}}[Z]\leq\mu\}.$
There is no single dominating measure for ${\mathcal{P}}$ since it contains the uncountable non-dominated family $\{\delta_{z}\colon z\in[0,\mu]\}$. However, ${\mathcal{P}}$ is generated by ${\mathcal{E}}_{0}=\{Z/\mu\}$ in the sense of Corollary 4.2, so our strategy will be to locate a candidate numeraire and then apply the corollary to verify that the candidate is, in fact, the numeraire. The alternative hypothesis ${\mathsf{Q}}$ is the uniform distribution on $[0,1]$, and we have ${\mathsf{Q}}\ll{\mathcal{P}}$ for the simple reason that ${\mathsf{P}}(A)=0$ for all ${\mathsf{P}}\in{\mathcal{P}}$ implies that $A$ must actually be empty.
To find a candidate numeraire, we observe that two natural e-variables are $Z/\mu$ and the constant one. All convex combinations of these are also e-variables; equivalently, $1+\lambda(Z-\mu)$ is an e-variable for each $\lambda\in[0,\mu^{-1}]$. We now look for a log-optimal e-variable in this class by directly maximizing $f(\lambda)={\mathbb{E}}_{\mathsf{Q}}[\log(1+\lambda(Z-\mu))]$ over $\lambda\in[0,\mu^{-1}]$. This is a strictly concave function whose derivative $f^{\prime}(\lambda)={\mathbb{E}}_{\mathsf{Q}}[(Z-\mu)/(1+\lambda(Z-\mu))]$ satisfies $f^{\prime}(0)=\frac{1}{2}-\mu>0$ and $f^{\prime}(\mu^{-1})={\mathbb{E}}_{\mathsf{Q}}[\mu-\mu^{2}/Z]=-\infty$. Thus there is a unique interior maximizer $\lambda^{*}\in(0,\mu^{-1})$ which is characterized by the first-order condition
${\mathbb{E}}_{\mathsf{Q}}\left[\frac{Z-\mu}{1+\lambda^{*}(Z-\mu)}\right]=0.$ (5.2)
Since ${\mathsf{Q}}$ is the standard uniform distribution we can be more explicit. Nothing changes if we first multiply both sides by $\lambda^{*}$, and then the left-hand side becomes
@@
${\mathbb{E}}_{\mathsf{Q}}\left[\frac{X}{X^{*}}\right]=1+(\lambda-\lambda^{*}){\mathbb{E}}_{\mathsf{Q}}\left[\frac{Z-\mu}{1+\lambda^{*}(Z-\mu)}\right]=1.$ (5.3)
@@
It is clear that every such $X$ is an e-variable, but it is less obvious that every e-variable is of this form. Nonetheless, equipped with this knowledge the calculation in (5.3) would directly yield the numeraire property of $X^{*}$. The usefulness of Corollary 4.2 is that a complete description of ${\mathcal{E}}$ is not required; it is enough to check (5.3) for $X=1$ and for $X$ in the generating set ${\mathcal{E}}_{0}$.
@@ 5.3 Sub-Gaussian with nonpositive mean
${\mathcal{P}}=\left\{{\mathsf{P}}\in M_{1}\colon{\mathbb{E}}_{\mathsf{P}}[e^{\lambda Z-\lambda^{2}/2}]\leq 1\text{ for all }\lambda\in[0,\infty)\right\}.$
The ‘one-sided’ restriction $\lambda\in[0,\infty)$ implies that $Z$ has nonpositive (potentially nonzero and even infinite) mean under any ${\mathsf{P}}\in{\mathcal{P}}$. Indeed, monotone convergence and the definition of ${\mathcal{P}}$ yield ${\mathbb{E}}_{\mathsf{P}}[Z]=\lim_{\lambda\downarrow 0}{\mathbb{E}}_{\mathsf{P}}[(e^{\lambda Z}-1)/\lambda]\leq\lim_{\lambda\downarrow 0}(e^{\lambda^{2}/2}-1)/\lambda=0$. As in the previous example, ${\mathcal{P}}$ does not admit any dominating measure. It is generated by the family ${\mathcal{E}}_{0}=\{e^{\lambda Z-\lambda^{2}/2}\colon\lambda\in[0,\infty)\}$, so we will again look for a candidate numeraire and verify it using Corollary 4.2. We let the alternative hypothesis ${\mathsf{Q}}$ be normal with mean $\mu>0$ and unit variance. As in the previous example, and for the same reason, we have ${\mathsf{Q}}\ll{\mathcal{P}}$.
To find a candidate numeraire we maximize ${\mathbb{E}}_{\mathsf{Q}}[\log X]$ over $X\in{\mathcal{E}}_{0}$. That is, we maximize ${\mathbb{E}}_{\mathsf{Q}}[\lambda Z-\lambda^{2}/2]=\lambda\mu-\lambda^{2}/2$ over $\lambda\in[0,\infty)$. The maximizer is $\lambda^{*}=\mu$, which yields the candidate
@@
${\mathbb{E}}_{\mathsf{Q}}\left[\frac{e^{\lambda Z-\lambda^{2}/2}}{e^{\mu Z-\mu^{2}/2}}\right]={\mathbb{E}}_{\mathsf{Q}}\left[e^{(\lambda-\mu)(Z-\mu)-(\lambda-\mu)^{2}/2}\right]=1.$ (5.4)
@@
${\mathbb{E}}_{\mathsf{Q}}\left[\frac{X}{X^{*}}\right]\leq\int_{[0,\infty)}{\mathbb{E}}_{\mathsf{Q}}\left[\frac{e^{\lambda Z-\lambda^{2}/2}}{e^{\mu Z-\mu^{2}/2}}\right]\pi(d\lambda)=1.$
@@ 5.4 Symmetric distributions
${\mathcal{P}}=\left\{{\mathsf{P}}\in M_{1}\colon Z\text{ and }-Z\text{ have the same distribution under }{\mathsf{P}}\right\},$
which is again a non-dominated family. We also fix an alternative hypothesis ${\mathsf{Q}}$ that admits a Lebesgue density $q$. It is natural to conjecture that the RIPr is given by the symmetrization $\widetilde{\mathsf{P}}$ of ${\mathsf{Q}}$, whose density is $\widetilde{p}(z)=(q(z)+q(-z))/2$. However, this cannot quite be true in general because $\widetilde{\mathsf{P}}$ need not be equivalent to ${\mathsf{Q}}$. Instead, we claim that the RIPr is the measure ${\mathsf{P}}^{*}$ with density
@@
This is the absolutely continuous part of $\widetilde{\mathsf{P}}$ with respect to ${\mathsf{Q}}$. It is a probability measure if ${\mathsf{Q}}$ has symmetric support, and otherwise a proper sub-probability measure. Note that $\widetilde{\mathsf{P}}$ belongs to ${\mathcal{P}}$, which in particular shows that ${\mathsf{Q}}\ll{\mathcal{P}}$ since ${\mathsf{Q}}\ll\widetilde{\mathsf{P}}$. (Alternatively, we again have that ${\mathsf{P}}(A)=0$ for all ${\mathsf{P}}\in{\mathcal{P}}$ implies that $A$ is empty, which also yields ${\mathsf{Q}}\ll{\mathcal{P}}$.)
To check that ${\mathsf{P}}^{*}$ is the RIPr we will show that the implied candidate numeraire is, in fact, the numeraire. It is given by
$X^{*}=\frac{d{\mathsf{Q}}}{d{\mathsf{P}}^{*}}=\frac{2q(Z)}{q(Z)+q(-Z)}.$
@@
where we recall that a function $\phi$ is odd if $\phi(z)+\phi(-z)=0$ for all $z\in{\mathbb{R}}$. Indeed, any $X$ of this form satisfies ${\mathbb{E}}_{\mathsf{P}}[X]\leq 1+{\mathbb{E}}_{\mathsf{P}}[\phi(Z)]=1$, where the symmetry of $Z$ and the oddness of $\phi$ were used to get ${\mathbb{E}}_{\mathsf{P}}[\phi(Z)]={\mathbb{E}}_{\mathsf{P}}[\phi(-Z)]=-{\mathbb{E}}_{\mathsf{P}}[\phi(Z)]$ and hence ${\mathbb{E}}_{\mathsf{P}}[\phi(Z)]=0$. Conversely, for any e-variable $X=f(Z)$ we may write $X=\frac{1}{2}(f(Z)+f(-Z))+\phi(Z)$ where $\phi(z)=\frac{1}{2}(f(z)-f(-z))$ is the odd part of $f(z)$. For any $z\in{\mathbb{R}}$ we use the symmetric distribution ${\mathsf{P}}=\frac{1}{2}(\delta_{z}+\delta_{-z})$ and the fact that $X$ is an e-variable to get $\frac{1}{2}(f(z)+f(-z))={\mathbb{E}}_{\mathsf{P}}[f(Z)]\leq 1$. Hence $X\leq 1+\phi(Z)$ as required. We could also have reached this conclusion by using that every admissible e-variable is of the form $1+\phi(Z)$ for some odd function $\phi$ as shown by Ramdas et al. (2020).
We can now verify the numeraire property. First, $X^{*}$ is ${\mathsf{Q}}$-almost surely strictly positive and finite, and it is an e-variable because it can be written as $X^{*}=1+\phi^{*}(Z)$ where
@@
${\mathbb{E}}_{\mathsf{Q}}\left[\frac{X}{X^{*}}\right]\leq{\mathbb{E}}_{\mathsf{Q}}\left[(1+\phi(Z))\frac{q(Z)+q(-Z)}{2q(Z)}\right]={\mathbb{E}}_{\widetilde{\mathsf{P}}}\left[(1+\phi(Z))\mathbf{1}_{\{q(Z)>0\}}\right]\leq{\mathbb{E}}_{\widetilde{\mathsf{P}}}[1+\phi(Z)]=1.$
This confirms the numeraire property. Finally, let us remark that it is not really necessary that ${\mathsf{Q}}$ admit a density. The symmetrization of ${\mathsf{Q}}$ is still well-defined as $\widetilde{\mathsf{P}}=\frac{1}{2}({\mathsf{Q}}+\widetilde{\mathsf{Q}})$ where $\widetilde{\mathsf{Q}}$ is the distribution of $-Z$ under ${\mathsf{Q}}$, or equivalently, the pushforward of ${\mathsf{Q}}$ under the reflection map $z\mapsto-z$. The RIPr is then the absolutely continuous part ${\mathsf{P}}^{*}=\widetilde{\mathsf{P}}^{a}$, and the numeraire is any nonnegative version of the Radon–Nikodym derivative $d{\mathsf{Q}}/d{\mathsf{P}}^{*}$ such that $d{\mathsf{Q}}/d{\mathsf{P}}^{*}-1$ is odd.
5.5 Exponential family with one-dimensional sufficient statistic
Here ${\mathcal{Z}}$ can be general and is equipped with a reference measure $\mu$. Consider an exponential family of densities
$p_{\theta}(z)=e^{\theta T(z)-A(\theta)}$
with respect to $\mu$, where $A$ is convex and differentiable, the sufficient statistic $T(z)$ is one-dimensional, and the natural parameter $\theta$ ranges in some interval $\Theta\subset{\mathbb{R}}$. The null hypothesis is ${\mathcal{P}}=\{p_{\theta}\colon\theta\in\Theta_{0}\}$ for some closed subset $\Theta_{0}\subset\Theta$, and the alternative is $q=p_{\theta_{1}}$ for some $\theta_{1}\in\Theta$. We suppose that $\Theta_{0}$ has a smallest element $\theta^{*}$ and that $\theta_{1}<\theta^{*}$. It is natural to conjecture that $p_{\theta^{*}}$ is the RIPr. To confirm this, it suffices to verify (4.3). Using the standard formula for the moment generating function of the sufficient statistic we get
${\mathbb{E}}_{\mathsf{Q}}\left[\frac{p_{\theta}}{p_{\theta^{*}}}\right]={\mathbb{E}}_{\mathsf{Q}}\left[e^{(\theta-\theta^{*})T(Z)-A(\theta)+A(\theta^{*})}\right]=e^{A(\theta_{1}+\theta-\theta^{*})-A(\theta_{1})-A(\theta)+A(\theta^{*})}$
for all $\theta\in\Theta_{0}$. Since $A$ is convex, its derivative $A^{\prime}$ is increasing. Together with the fundamental theorem of calculus, as well as $\theta_{1}<\theta^{*}$ and $\theta-\theta^{*}\geq 0$ for $\theta\in\Theta_{0}$, this gives
[…]
$=A(\theta)-A(\theta^{*}).$
We conclude that ${\mathbb{E}}_{\mathsf{Q}}[p_{\theta}/p_{\theta^{*}}]\leq 1$ for all $\theta\in\Theta_{0}$ so that (4.3) holds and $p_{\theta^{*}}$ is indeed the RIPr. As a result the numeraire is the likelihood ratio
$X^{*}=\frac{p_{\theta_{1}}(Z)}{p_{\theta^{*}}(Z)}=e^{(\theta_{1}-\theta^{*})T(Z)-A(\theta_{1})+A(\theta^{*})}.$
We refer to Csiszar and Matus (2003) for an extensive study of information projections involving exponential families.
Remark 5.3.
When ${\mathcal{P}}$ is a parametric family of distributions it is often easier to first find the RIPr and then use it to find the numeraire. Here we used Theorem 4.7 to confirm that the candidate $p_{\theta^{*}}$ was indeed the RIPr. We could also have minimized relative entropy and used Corollary 4.10. Since the corollary assumes that ${\mathcal{P}}$ is convex, which fails here, we would have had to apply it with the convex hull of ${\mathcal{P}}$ and perform the minimization over this set.